A solution contains 20g of solute in 180g of solvent. If the solvent is water, what is the concentration of the solution in terms of mass by mass percentage



A. 11.%


B. 22.2%


C. 10%


D. 20%


Please show working

1
Expert's answer
2022-02-01T02:49:02-0500

Concentration of a solution:

"m/m\space=\frac{mass\space of \space solute}{mass\space of\space solution}×100"

Mass of solutions=mass of solute + mass of solvent =20g+180g = 200g

"m/m\space=\frac{20}{200}×100=10%"